OpenStudy (anonymous):

Manny loves to eat Cajun pizza. Each square inch of pizza is 8 calories. He orders a slice of pizza from a pizza with a diameter of 16 inches that is cut with a central angle of 20 degrees. How many calories does he eat? about 89 calories about 11 calories about 1,609 calories about 201 calories

OpenStudy (jewlzme17):

20/360 = x / (pi 8^2) where x is the area of the pizza slice 1/18 = x/64 pi 18x = 64 pi x = 11.170 square inches At 8 calories per square inch, the person consumes 8 * 11.170 calories = 89.360 calories (approx)